# Diffusive contact between randomly driven colloidal suspensions

### v1

@@ Abstract
We study the relaxation process of two driven colloidal suspensions in diffusive contact to a steady state, similar to thermalization. We start by studying a single suspension, subjecting it to random driving forces via holographic optical tweezers, which agitate it to a higher effective temperature. Interestingly, the effective temperature of the suspension, defined by the Einstein relation, exhibits a non-monotonic dependence on the driving frequency. Next, we follow the flux of particles between two such suspensions in diffusive contact, starting from a uniform density and relaxing to a state with zero net particle flux. The density remains uniform for systems with different frequencies but equal effective temperatures. At high driving frequencies, we show that the density distribution at steady state is determined by equating the ratio of the chemical potential to the effective temperature in both systems, mirroring thermal equilibrium behavior.
@@
A dilute suspension of spherical active or randomly driven colloidal particles provides a unique platform for investigating non-equilibrium statistical physics. Although clearly in a non-equilibrium state, this system often exhibits well-defined state functions. First, the osmotic pressure, denoted as $\Pi$, coincides with the thermodynamic pressure when defined mechanically, under the condition that interactions between colloidal particles and between them and the system’s boundaries are torque-free [1, 2, 3, 4]. Second, an effective temperature, denoted as $T_{\text{eff}}$, is well-defined through the generalized fluctuation-dissipation relation (GFDR) when active collisions occur at shorter timescales than particles’ thermal relaxation to the potential landscape’s minima. [5, 6, 7, 8].
@@
Notably, these suspensions adhere to an ideal-gas-like equation of state similar to the equation of state observed in dilute thermal colloidal suspensions $\Pi=nk_{B}T$,[9, 10, 11], with $T\to T_{\text{eff}}$, where $n$ is the particle number density and $k_{B}$ the Boltzmann constant [12, 13, 1]. Experiments [13] and simulations [14, 1] have explored deviations from this equation of state as colloidal density increases. These timescales could arise from factors like the reorganization of the suspension or memory effects from complex collision dynamics [15].
The concept of effective temperature, a macroscopic variable governing heat flow independent of microscopic details, has proven valuable in non-equilibrium thermodynamics [16, 17]. Despite its widespread use, a crucial gap exists in the literature regarding the quantification of the effective temperature’s sensitivity to microscopic details. This information is essential for defining its applicability and limitations. Furthermore, identifying state functions that predict equilibration upon contact between driven or active systems would offer significant insights into the underlying non-equilibrium thermodynamics governing these materials.
@@
Here, we address both topics, creating systems of driven colloidal suspensions, allowing us to change the microscopic dynamics of the particles in a controlled manner. To this aim, we apply random optical forces to the colloidal particles and measure the resulting suspension’s effective temperature. We find that we can obtain the same effective temperature with two different driving protocols inducing different microscopic dynamics. When well defined, the effective temperature governs the thermodynamics of the driven suspensions independently of the underlying microscopic dynamics. Next, using the ideal-gas-like description of colloidal suspensions, we show that the equilibration between driven colloidal suspension in diffusive contact is reached when the ratio of the chemical potential and the effective temperature is equal on both sides, in accord with equilibrium thermodynamics.
We used holographic optical tweezers (HOTs) [18, 19, 20] to generate multiple randomly placed optical beams (Fig. 1) and switch their position at constant time intervals $\tau_{s}=1/f$. Our HOTs setup, previously described in detail [21], is based on a Keopsys fiber laser (wavelength, $1083$ nm). We control the position of the beams by imprinting a phase-only computer-generated hologram on the laser beam with a spatial light modulator (LCOS-SLM, X10468, Hamamatsu) capable of reliably switching its pattern with a frequency $f\leq 20$ Hz. We couple the laser pattern on the SLM to the back aperture of a 100x oil immersion objective (NA=1.42) mounted on an Olympus IX71 inverted microscope. The light pattern is then formed in the sample plane.
@@
We drive the colloidal suspension, mimicking thermal forces in the following manner. An array of 36 randomly positioned beams is projected onto a region of interest (ROI) in the sample plane. To ensure homogeneous driving, we divide the ROI into a fictitious square array with a lattice parameter of $5.0~{}\pm$ $0.1~{}\mu m$. Each laser beam is assigned to one square in the lattice, i.e., each beam location is generated randomly within its respective square (Fig 1). The random distance between the laser beams and colloidal particles diffusing in the suspension ensures that particles are subjected to forces of varying strengths and directions. However, unlike thermal white noise, these tailored fluctuating forces change at constant time intervals.
@@
The resulting motion of the colloidal particles is a type of Run-and-Tumble motion, where the direction of motion switches completely between runs with no memory or preference for directions. During the runs, the particle is pushed away from the laser beam towards the minimum of the potential landscape. The motion of a single particle of position $x(t)$ under the influence of a laser beam located at $y(t)$ can be described using a modified Langevin equation similar to that of a self-propelling particle [3],
$\frac{d\vec{x}(t)}{dt}=\frac{\vec{F}_{\text{opt}}(t)}{\gamma}+\sqrt{2D}\vec{\eta}(t),$
$\vec{F}_{\text{opt}}=\frac{A}{\sigma^{2}}(\vec{x}-\vec{y})e^{\frac{-(\vec{x}-\vec{y})^{2}}{2\sigma^{2}}},$ (1)
where we have neglected inter-particle interactions. Here, $\vec{F}_{\text{opt}}$ is the stochastic Gaussian-shaped optical force driving the particle, with beam amplitude $A$ and width $\sigma$. The beam’s random position $y$ changes at a frequency $f$, $D$ is the diffusion constant, and $\eta$ is a Gaussian white noise of unit variance.
@@
Figure 2: Extracting the effective temperature, $T_{D}$, from particle trajectories. a) Probability distribution of particle displacement, at $f=1$ Hz, with lag time $\tau=1/30,1/10,1$ s (symbols) fitted to a Gaussian distribution (solid line). Inset: a typical trajectory of a particle at $f=1$ Hz. b) $\langle\Delta r^{2}\rangle$ as a function of $\tau$, for experimental measurements of $f=0.1,1,10,20$ Hz. c) Experimental measurements (filled diamonds) and simulation results (empty squares) of the diffusion coefficient, friction coefficient, and effective temperature normalized by their value in the absence of driving as a function of switching frequency $f$.
@@
To characterize the relation between the driving frequency and the resulting effective temperature of the colloidal suspension, we perform a set of experiments in which we change the driving frequency in the range of $0.1~{}\text{Hz}\leq f\leq 20~{}\text{Hz}$. We record 10 movies of 5.5 min at 30 fps using a CMOS camera (GS3-U3-2356M, FLIR) for each driving frequency from which we extract particle trajectories using standard particle tracking algorithms [22], see, for example, a typical trajectory at a driving frequency of $f=1$ Hz (Fig. 2a, inset) [23]. The position probability distribution function (PDF) and mean square displacement $\langle\Delta r^{2}\rangle$ of the suspended particles are then calculated (Fig. 2a,b). Since the PDF is Gaussian at $\tau>1$ s and $\langle\Delta r^{2}\rangle$ increases linearly with lag time, $\tau$, we treat the long-time particle motion as normal diffusion [24]. The diffusion coefficient is then given by $D=\langle\Delta r^{2}\rangle/4\tau$.
@@
In Fig. 2c (top pannel), we plot $D/D_{0}$ as a function of the driving frequency, where $D_{0}$ is the diffusion coefficient at room temperature with no driving. Interestingly, $D$ is non-monotonic in $f$, reaching a maximum value at $f\sim 2$ Hz. This behavior can be understood by looking at the two limiting cases. At the limit of very slow driving, the projected pattern barely changes while particles diffuse freely, avoiding the laser beams. Thus, we expect $D/D_{0}\simeq 1$. In contrast, at very fast driving, the light pattern changes faster than the relaxation of the particles to the new potential landscape. Therefore, the particles experience a time-averaged uniform potential in which they diffuse freely, resulting in $D/D_{0}\simeq 1$. We expect $D$ to be maximal when the relaxation time of the particles to the new minima of the potential landscape, $\tau_{r}$, is equivalent to the switching time $\tau_{s}$ of the pattern. In this regime, the particle’s response to the optical forces results in maximal motion, leading to higher diffusivity levels and a higher effective temperature.
@@
The friction coefficient of the suspension in the presence of driving is measured from a second set of experiments, in which we translate the suspension slowly through the light pattern and measure the average particle velocity [25]. The friction coefficient is given by $\gamma=\gamma_{0}v_{0}/v$, where $\gamma_{0}$ and $v_{0}$ represent the friction coefficient and average velocity of the suspended particles without external driving. Figure 2c (middle pannel) shows the weak dependence of $\gamma$ on the driving frequency.
@@
Following previous studies [12, 5, 6, 3, 26], we use the Einstein relation to define the effective temperature of the suspension, $k_{B}T_{D}=\gamma D$. The effective temperature shown in Fig. 2c (bottom panel), exhibits a non-monotonic dependence on frequency. This behavior mirrors the trend observed in the diffusion coefficient, as the friction coefficient remains relatively constant across the measured frequency range.
@@
We complement our experiments with Brownian dynamics simulation to extend the range of frequencies and driving strengths examined. In our simulations, particles are subjected to thermal and optical forces implemented according to Eq. 1. Particles interact via hard core repulsion implemented by the WCA potential [27]. In the first set of simulations, we use $A=35k_{B}T$, $T=293$ K, $\gamma_{0}=6.178\cdot 10^{-8}$ kg/s, and $\sigma=0.64~{}\mu$m similar to the experimental conditions, and ran 10 simulations of 10 min duration to extract $D$ and $\gamma$ as a function of $f$ (Fig. 2c). We observe only a qualitative agreement between simulations and experiments, mainly due to the lack of hydrodynamic interactions. However, simulations show a maximum effective temperature at a similar switching frequency as experiments (Fig. 2c) [28].
@@
Figure 3: Driving characteristics extracted form simulated data with $A=35k_{B}T$. A typical optical force acting on a single particle as a function of time for a) $f=0.1$ Hz, and b) $f=50$ Hz. c) The persistence particle motion time $\tau_{p}$ as a function of the switching frequency. d) The average force acting on the particle as a function of the switching frequency.
@@
We conducted further simulations in which we significantly enhanced the repulsion force of the laser beams (6 fold) and increased the driving frequency beyond experimentally attainable values (up to 200 Hz). This extended the range of examined conditions and allowed us to enhance the effect of optical forces in relation to thermal motion [7, 8].
@@
From the simulated data, we can readily calculate the force acting on each colloidal particle as a function of time, as depicted for $f=0.1$ Hz and $f=50$ Hz in Fig. 3a,b, respectively. The initial amplitude of the force is determined by the random distance between the closest beam to each particle. The persistence time of particle motion is given by $\tau_{p}=\text{min}(\tau_{s},\tau_{r})$. As the driving frequency increases, $\tau_{p}$ transitions from $\tau_{r}$, 0.72 s in our simulations, to $\tau_{s}$ (Fig. 3c). The average force amplitude acting on the particles increases with frequency and saturates towards a finite value (Fig. 3d [29]). For Run-and-Tumble particles $k_{B}(T_{D}-T)=v^{2}\gamma/2\alpha$ where $v$ is the run speed and $\alpha$ is the tumble rate for $t\gg\alpha^{-1}$ [3]. By substituting $v=F_{\text{av}}\gamma$ and $\alpha^{-1}=\tau_{p}$ we obtain $k_{B}(T_{D}-T)\sim F_{\text{av}}^{2}\tau_{p}/\gamma$. The coupled effect of the non-linear change in mean force amplitude and persistent time further justifies the non-monotonic change in the effective temperature observed in our experiments and simulations.
@@
Figure 4: Experimental measurement of particle flow between two driven systems in diffusive contact as illustrated on the right panel. The probability of finding a particle on the side of the higher effective temperature is plotted as a function of time for three typical cases: (1) same driving frequency and same effective temperature (orange), (2) different driving frequency but same effective temperature (purple), and (3) different driving frequency and different effective temperatures (green). Net particle flow is observed only when the effective temperature is different on the two sides. Here, $f_{1}=1$ Hz, $f_{2}=2$ Hz, $f_{3}=0.1$ Hz resulting in $T_{1}/T=T_{2}/T=1.7\pm 0.2$ and $T_{3}/T=1\pm 0.2$.
@@
In equilibrium, the transport of particles between two systems in diffusive contact is determined by the temperature and chemical potential [30]. The conventional assumption that thermalization is achieved by maximizing entropy with respect to energy and particle flow leads to two requirements, $T_{1}=T_{2}$ and $\mu_{1}/T_{1}=\mu_{2}/T_{2}$, respectively. The chemical potential of dilute colloidal suspensions is given by [31], $\mu=-k_{B}T\ln{[(\rho_{0}/\rho)(T/T_{0})^{3/2}]}$, where $\rho$ is the number density of colloidal particles and $\rho_{0},T_{0}$ refer to some reference state. Following the generalization of pressure in terms of the effective temperature, we generalize the chemical potential of our randomly driven suspensions by substituting $T\to T_{D}$.
@@
With this analogy in mind, our objective is to investigate the particle flow between two regions with different effective temperatures over time. To achieve this, we divide our ROI into two subsystems, each driven independently with a different switching frequency, as depicted in the right panel of Fig. 4. All experiments start with a uniform distribution of particles. We measure the probability of observing a particle in the system at the higher effective temperature $P_{\text{hot}}$ as a function of time. In our control experiment, we drive both sides of the system independently but with the same frequency $f_{1}=1$ Hz, resulting in zero average current (orange line in Fig 4). Next, we chose a different frequency $f_{2}=2$ Hz that drives the suspension to a similar effective temperature $T_{1Hz}/T\sim T_{2Hz}/T=1.7\pm 0.2$. We use $f_{1}$ as the driving frequency on one side of the system and $f_{2}$ on the other. While the effective temperature is similar on both sides, particle dynamics are different. Nonetheless, we observe no net particle flow between the two sides (purple line Fig 4). This result indicates that the effective temperature used here may be a state function of the system and is independent of microscopic dynamics. Finally, we drive the two sides of the system with different frequencies, $f_{1}=1$ Hz and $f_{3}=0.1$ Hz, resulting in different effective temperatures, $T_{1}/T=1.7\pm 0.2$ and $T_{3}/T=1\pm 0.2$. Under these conditions particles flow towards the side with a lower effective temperature (green line Fig. 4 [32]).
@@
Since the effective temperature is kept constant by our driving protocol, and particles can not carry with them effective heat, we expect that the two systems in contact will reach a steady state in which $\mu/k_{B}T_{D}$ is equal on both sides. This condition is held trivially in our first two experiments since the effective temperature $T_{D}$ and particle density $\rho$ are equal from the start. However, for the third experiment, we have $\mu_{1}/k_{B}T_{3}\sim-14.7\pm 0.8$ and $\mu_{3}/k_{B}T_{3}\sim-13.3\pm 1.3$. These values are equal within experimental measurement, suggesting that in dilute colloidal suspension, the effective temperature defined through the Einstein relation determines the thermodynamic behavior of the suspensions.
@@
Figure 5: Comparison of the thermodynamic variables of two driven suspensions in diffusive contact (squares and circles) (a) The chemical potential as a function of the minimal switching frequency $f_{\text{min}}$. (b) The effective temperature as a function of $f_{\text{min}}$. (c) $\mu/k_{B}T$ as a function of $f_{\text{min}}$. Simulations were run with $A=200k_{B}T$.
@@
Next, we conduct a series of simulations to investigate the spectrum of driving frequencies wherein equilibrium conditions for reaching a steady state apply. Given that the GFDR holds for $\tau_{s}<\tau_{r}$, we examine our findings through a plot depicting the lower driving frequency, $f_{\text{min}}$, of the two interacting systems. Specifically, we analyze both systems’ effective chemical potential, temperature, and ratio (Fig. 5). We find that as $f_{\text{min}}$ increases, the values of $\mu$ and $T_{D}$ change less with the driving frequency. Moreover, while the values of $\mu/T_{D}$ of both sides agree within measurement error at steady-state, the agreement is very good when $f_{\text{min}}>1/\tau_{r}$. Hence, we conclude that within our driving scheme, at sufficiently high frequencies, the driven colloidal suspensions behave as equilibrium colloidal suspensions at elevated temperatures.

### v2

@@ Abstract
We study the relaxation process of two driven colloidal suspensions in contact, to a joint steady state, similar to the process of thermalization. First, we study a single suspension, subjecting it to random driving forces via holographic optical tweezers, which agitate it to a higher effective temperature. Interestingly, the effective temperature of the suspension, defined by the Einstein relation, exhibits a non-monotonic dependence on the driving frequency. Next, we follow the flux of particles between two such suspensions in diffusive contact, starting from a uniform density and relaxing to a state with zero net particle flux. At high driving frequencies, we show that the density distribution at steady state is determined by equating the ratio of the chemical potential to the effective temperature in both systems, reminiscent of the thermal equilibrium behavior.
American Chemical Society,
@@
A dilute suspension of spherical, randomly driven colloidal particles provides a unique platform for investigating non-equilibrium statistical physics. Although clearly in a non-equilibrium state, such systems often exhibit well-defined state functions. First, the osmotic pressure, denoted as $\Pi$, coincides with the thermodynamic pressure when defined mechanically, under the condition that interactions between colloidal particles and between them and the boundaries of the system are torque-free [1, 2, 3, 4, 5]. Second, an effective temperature, denoted as $T_{\text{eff}}$, is well-defined through the generalized fluctuation-dissipation relation (GFDR) when active collisions occur at shorter timescales than the thermal relaxation of the particles to the potential landscape’s minima. [6, 7, 8, 9].
@@
Notably, these suspensions exhibit an equation of state analogous to that of ideal gases, mirroring the behavior observed in dilute thermal colloidal suspensions: $\Pi=nk_{B}T_{\text{eff}}$. Here, $n$ represents the particle number density, $k_{B}$ is the Boltzmann constant, and $T_{\text{eff}}$ is an effective temperature that replaces the thermal temperature $T$ in the traditional equation [10, 11, 12]. Experiments [13] and simulations [14, 1] have explored deviations from this equation of state as colloidal density increases. Such deviations could arise from the reorganization of the suspension or from memory effects due to complex collision dynamics [15].
Despite the widespread use of effective temperatures [16, 17], a significant gap remains in the literature concerning the role of effective temperatures in the relaxation process between two systems in contact as they evolve toward a steady state. For instance, thermal equilibration between systems is independent of their nature: tea cools in air and can also cool by contact with ice. In contrast, far from equilibrium, the effective temperature is often defined in a system-specific manner, rendering the extension of the notion of heat flow between different systems in contact ill-defined. Identifying state functions, such as effective temperature or effective chemical potential, that predict equilibration between driven or active systems in contact is crucial for the generalization of thermodynamics to non-equilibrium conditions.
@@
In this work, we investigate the role of GFDR-based effective temperature in the relaxation to a driven steady state of two driven colloidal suspensions in diffusive contact. Our setup allows us to disentangle the influence of driving dynamics from that of the resulting effective temperature on the equilibration process. Specifically, we create a diffusive contact between two driven systems and measure the particle flow between them. We define an effective chemical potential that depends on the effective temperature and demonstrate that no particle flow occurs when the two systems share the same effective temperature and effective chemical potential, even when driven by different driving protocols. Additionally, we show that the ratio of the generalized effective chemical potential and the effective temperature governs the steady-state particle density in each system when the effective temperature in the two sides is different.
To drive colloidal suspensions in a controlled manner, we apply random optical forces to the colloidal particles and measure the resulting diffusion coefficient $D$ and drag $\gamma$. We calculate the effective temperature using a generalization of the Einstein relation, $T_{\text{{eff}}}=\gamma D$. We find that the effective temperature is non-monotonic with driving frequency. Specifically, two different driving frequencies can result in the same effective temperature, allowing us to examine the effective temperature’s sensitivity to microscopic kinetic details. We find that when well defined, the effective temperature governs the thermodynamics of the driven suspensions independently of the underlying microscopic dynamics. Next, using the ideal-gas-like description of colloidal suspensions, we define an effective chemical potential and show that the equilibration between driven colloidal suspensions in diffusive contact is reached when the ratio of the chemical potential and the effective temperature is equal on both sides, in accord with the entropy maximization rule.
We used holographic optical tweezers (HOTs) [18, 19, 20] to generate multiple randomly placed optical beams (Fig. 1) and switch their position at constant time intervals $\tau_{s}=1/f$. Our HOTs setup, previously described in detail [21], is based on a Keopsys fiber laser (wavelength, $1083$ nm). We control the position of the beams by imprinting a phase-only computer-generated hologram on the laser beam using a spatial light modulator (LCOS-SLM, X10468, Hamamatsu) capable of reliably switching its pattern with a frequency $f\leq 20$ Hz. We couple the laser pattern on the SLM to the back aperture of a 100x oil immersion objective (NA=1.42) mounted on an Olympus IX71 inverted microscope. The light pattern is then formed in the sample plane.
@@
We drive the colloidal suspension, mimicking thermal forces in the following manner. An array of 36 randomly positioned beams is projected onto a region of interest (ROI) in the sample plane. To ensure homogeneous driving, we divide the ROI into a fictitious square array with a lattice parameter of $5.0~{}\pm$ $0.1~{}\mu m$. Each laser beam is assigned to one square in the lattice, i.e., each beam location is generated randomly within its respective square (Fig. 1). The random distance between the laser beams and colloidal particles diffusing in the suspension ensures that particles are subjected to forces of varying strengths and directions. The dimensions of the fictitious square array were chosen to obtain a large region of driving without compromising the driving strength (see Fig. S1, Fig. S2 for driving force characterization). However, unlike thermal white noise, these tailored fluctuating forces change at constant time intervals.
@@
The resulting motion of the colloidal particles is a type of Run-and-Tumble motion, where the direction of motion switches completely between runs with no memory or preference for directions. During the runs, the particle is pushed away from the laser beam towards the minimum of the potential landscape by a Gaussian-shaped optical force $\vec{F}_{\text{opt}}=\frac{A}{\sigma^{2}}(\vec{x}-\vec{y})e^{\frac{-(\vec{x}-\vec{y})^{2}}{2\sigma^{2}}}$, where $A$ and $\sigma$ are the optical beam’s amplitude and width, respectively, $\vec{x}$ is the position of the particle and $\vec{y}$ is the position of the beam. The beam’s random position $\vec{y}$ changes randomly at a frequency $f$, and $\eta$ is a Gaussian white noise of unit variance.
@@
Figure 2: Extracting the effective temperature, $T_{\text{eff}}$, from particle trajectories. a) Probability distribution of particle displacement, at $f=1$ Hz, with lag time $\tau=1/30,1/10,1$ s (symbols) fitted to a Gaussian distribution (solid line). Inset: a typical trajectory of a particle at $f=1$ Hz. b) $\langle\Delta r^{2}\rangle$ as a function of $\tau$, for experimental measurements of $f=0.1,1,20$ Hz. c) Experimental measurements (filled diamonds) and simulation results (empty squares) of the diffusion coefficient, friction coefficient, and effective temperature normalized by their value in the absence of driving as a function of switching frequency $f$, where $T_{0}\approx$ 293 K is the measured ambient temperature, $D_{0}=$ 0.065 $\mu\text{m}^{2}/\text{s}$ is the diffusion coefficient of a free particle, and $\gamma_{0}=6.18*10^{-8}$ kg/s is the drag coefficient calculated via the Einstein relation.
@@
To characterize the relation between the driving frequency and the resulting effective temperature of the colloidal suspension, we perform a set of experiments in which we change the driving frequency in the range of $0.1~{}\text{Hz}\leq f\leq 20~{}\text{Hz}$. We record 10 movies of 5.5 min at 30 fps using a CMOS camera (GS3-U3-2356M, FLIR) for each driving frequency from which we extract particle trajectories using standard particle tracking algorithms [22]. A typical trajectory at a driving frequency of $f=1$ Hz is shown in the inset of Fig. 2a and in Fig. S3. The probability distribution function (PDF) and mean square displacement $\langle\Delta r^{2}\rangle$ of the suspended particles are then calculated (Fig. 2a,b). Since the PDF is Gaussian at $\tau>1$ s and $\langle\Delta r^{2}\rangle$ increases linearly with lag time, $\tau$, we treat the long-time particle motion as normal diffusion (see Fig. S4 for more details, with Ref. [1]). The diffusion coefficient is then given by $D=\langle\Delta r^{2}\rangle/4\tau$.
@@
In Fig. 2c (top panel), we plot $D/D_{0}$ as a function of the driving frequency, where $D_{0}$ is the diffusion coefficient at room temperature with no driving. Interestingly, $D$ is non-monotonic in $f$, reaching a maximum value at $f\sim 2$ Hz. This behavior can be understood by looking at the two limiting cases. At the limit of very slow driving, the projected pattern barely changes while particles diffuse freely, avoiding the laser beams (see SM Movie 1). Thus, we expect $D/D_{0}\simeq 1$. In contrast, at very fast driving, the light pattern changes faster than the relaxation of the particles to the new potential landscape. Therefore, the particles experience a time-averaged uniform potential in which they diffuse freely, resulting in $D/D_{0}\simeq 1$ (see SM Movie 2). The maximum in $D$ emerges at an intermediate frequency where the timescale of pattern switching, $\tau_{s}$, matches the particle relaxation time, $\tau_{r}$. When $\tau_{s}\sim\tau_{r}$, the system is optimally tuned: particles can fully respond to each new configuration of the optical landscape before it changes again. This condition maximizes the particle response to the time-dependent potential: the particles are displaced significantly by each change in the landscape before it shifts again, while minimizing the time in which the particles are not affected by the optical beams per each cycle. This results in a peak in both diffusivity and effective temperature.
@@
We determine the friction coefficient $\gamma$ of the suspension in the presence of driving by a second set of experiments, in which we translate the suspension slowly through the light pattern. We repeat the experiments for each driving frequency, measuring the average particle velocity (see Fig. S5 for a detailed description of these experiments). The friction coefficient is given by $\gamma=\gamma_{0}v_{0}/v$, where $\gamma_{0}$ and $v_{0}$ represent the friction coefficient and average velocity of the suspended particles without external driving. Figure 2c (middle panel) shows the weak dependence of $\gamma$ on the driving frequency.
@@
Following previous studies [23, 6, 7, 3, 24], we use the Einstein relation to define the effective temperature of the suspension, $k_{B}T_{\text{eff}}=\gamma D$. The effective temperature shown in Fig. 2c (bottom panel) exhibits a non-monotonic dependence on frequency. This behavior mirrors the trend observed in the diffusion coefficient, as the friction coefficient remains relatively constant across the measured frequency range.
@@
We complement our experiments with Brownian dynamics simulations, numerically integrating the Langevin equation, to extend the range of frequencies and driving strengths examined. In our simulations, particles are subjected to thermal and optical forces. Particles interact via hard core repulsion implemented by the WCA potential. In the first set of simulations, we use $A=35~{}k_{B}T$, $T_{0}=293$ K, $\gamma_{0}=6.178\cdot 10^{-8}$ kg/s, and $\sigma=0.64~{}\mu$m, similar to the experimental conditions, and collect data from 10 simulations of 10 min duration each to extract $D$ and $\gamma$ as a function of $f$.
We observe a qualitative agreement between simulations and experiments (Fig. 2c). Namely, the diffusion coefficient and the effective temperature peak around $f=2$ Hz, while the friction coefficient barely changes with frequency. A slightly lower effective temperature is observed in the simulations, consistent with an enhancement of driving due to hydrodynamic interactions present in the experiments. This effect is minor since hydrodynamic interactions near the cell floor are suppressed (see Fig. S6 and discussion, which includes Refs.[2, 3, 4]).
@@
Figure 3: Driving characteristics extracted from simulated data with $A=35~{}k_{B}T$. An example of a typical combined force F, (optical and WCA) acting on a single particle as a function of time for a) $f=0.1$ Hz, and b) $f=50$ Hz. c) The persistence particle motion time $\tau_{p}$ as a function of the switching frequency. d) The time ${}_{\text{av}}$,
acting on the particle as a function of the switching frequency.
@@
We conducted further simulations in which we significantly enhanced the repulsion force of the laser beams (6 fold) and increased the driving frequency beyond experimentally attainable values (up to 60 Hz). This extended the range of examined conditions and allowed us to enhance the effect of optical forces in relation to thermal motion [8, 9].
@@
From the simulated data, we can readily calculate the force acting on each colloidal particle as a function of time, as depicted for $f=0.1$ Hz and $f=50$ Hz in Fig. 3a,b, respectively. The initial amplitude of the force is determined by the random distance between the closest beam to each particle. The persistence time of particle motion is given by $\tau_{p}=\text{min}(\tau_{s},\tau_{r})$. As the driving frequency increases, $\tau_{p}$ transitions from $\tau_{r}$, 0.72 s in our simulations, to $\tau_{s}$ (Fig. 3c). The average force amplitude acting on the particles increases with frequency and saturates towards a finite value (Fig. 3d, Fig. S7). For Run-and-Tumble particles $k_{B}(T_{\text{eff}}-T)=v^{2}\gamma/2\alpha$ where $v$ is the run speed and $\alpha$ is the tumble rate for $t\gg\alpha^{-1}$ [3]. By substituting $v=F_{\text{av}}\gamma$ and $\alpha^{-1}=\tau_{p}$ we obtain $k_{B}(T_{\text{eff}}-T)\sim F_{\text{av}}^{2}\tau_{p}/\gamma$. The coupled effect of the non-linear change in mean force amplitude and persistent time further justifies the non-monotonic change in the effective temperature observed in our experiments and simulations.
@@
Figure 4: Experimental measurement of particle flow between two driven systems in diffusive contact, as illustrated on the right panel. The probability of finding a particle on the region of the higher effective temperature is plotted as a function of time for three typical cases: (1) same driving frequency and same effective temperature (orange), (2) different driving frequency but same effective temperature (purple), and (3) different driving frequency and different effective temperatures (green). Net particle flow is observed only when the effective temperature is different on the two sides. Here, $f_{1}=1$ Hz, $f_{2}=2$ Hz, $f_{3}=0.1$ Hz resulting in $T_{1}/T_{0}=T_{2}/T_{0}=1.7\pm 0.2$ and $T_{3}/T_{0}=1.0\pm 0.2$.
@@
In equilibrium, the transport of particles between two systems in diffusive contact is determined by the temperature and chemical potential [25]. The conventional assumption that thermalization is achieved by maximizing entropy with respect to energy and particle flow leads to two requirements, $T_{1}=T_{2}$ and $\mu_{1}/T_{1}=\mu_{2}/T_{2}$, respectively. The chemical potential of dilute colloidal suspensions is given by [26],
$\mu=-k_{B}T\ln{[(\rho_{0}/\rho)(T/T_{0})^{3/2}]},$ (1)
where $\rho$ is the number density of colloidal particles at temperature $T$, and $\rho_{0}$ and $T_{0}$ refer to a reference state. Following the generalization of osmotic pressure in terms of the effective temperature, we generalize the chemical potential of our randomly driven suspensions by substituting $T\to T_{\text{eff}}$ to obtain $\mu_{\text{eff}}\equiv\mu(T_{\text{eff}})$.
@@
With this analogy in mind, our objective is to investigate the particle flow between two regions with different effective temperatures over time. To achieve this, we divide our ROI into two subsystems, each driven independently with a different switching frequency, as depicted in the right panel of Fig. 4 and SM movie 3. All experiments start with a uniform distribution of particles. In the diffusive contact experiment, we generated an optical “cage” to ensure that no particles enter or exit the area of the two subsystems. The cage consists of 56 beams, 1.38 microns apart, forming a square (see SM Movie 4). The addition of optical beams results in weaker driving beams than the previous set of experiments. The effective temperature, which depends linearly on the driving force, is, therefore, smaller in these experiments. We measure the probability of observing a particle in the system at the higher effective temperature region, $P_{\text{hot}}$, as a function of time. In our control experiment, we drive both sides of the system independently but with the same frequency $f_{1}=1$ Hz, resulting in zero average current (orange line in Fig. 4). Next, we choose a different frequency $f_{2}=2$ Hz that drives the suspension to a similar effective temperature $T_{1}/T_{0}\sim T_{2}/T_{0}=1.7\pm 0.2$. This choice ensures a well-defined effective temperature on both sides [8]. We use $f_{1}$ as the driving frequency on one side of the system and $f_{2}$ on the other. While the effective temperature is similar on both sides, the particle dynamics are different. Nonetheless, we observe no net particle flow between the two sides (purple line Fig 4). This result indicates that the effective temperature used here may be a state function of the system and is independent of microscopic dynamics. Finally, we drive the two sides of the system with different frequencies, $f_{1}=1$ Hz and $f_{3}=0.1$ Hz, resulting in different effective temperatures, $T_{1}/T_{0}=1.7\pm 0.2$ and $T_{3}/T_{0}=1.0\pm 0.2$. Under these conditions particles flow toward the side with a lower effective temperature (green line Fig. 4). We further demonstrate through simulations that this mechanism is reversible when the conditions in the two contacting baths are switched back and forth (see Fig. S8).
@@
Since the effective temperature is kept constant by our driving protocol, and particles cannot carry with them effective heat, we expect that the two systems in contact will reach a steady state in which $\mu_{\text{eff}}/k_{B}T_{\text{eff}}$ is equal on both sides. This condition is held trivially in our first two experiments since the effective temperature $T_{\text{eff}}$ and particle density $\rho$ are equal from the start. However, for the third experiment, we have $\mu_{1}/k_{B}T_{3}\sim-14.7\pm 0.8$ and $\mu_{3}/k_{B}T_{3}\sim-13.3\pm 1.3$. These values are equal within experimental measurement, suggesting that in dilute colloidal suspension, the effective temperature defined through the Einstein relation determines the thermodynamic behavior of the suspensions.
@@
Figure 5: Comparison of the thermodynamic variables of two driven suspensions in diffusive contact (empty and filled orange squares for simulations and blue diamonds for experiments). (a) The chemical potential as a function of the minimal switching frequency $f_{\text{min}}$. (b) The effective temperature as a function of $f_{\text{min}}$. (c) $\mu_{\text{eff}}/k_{B}T_{\text{eff}}$ as a function of $f_{\text{min}}$. Simulations were run with $A=200k_{B}T$. Error bars were estimated using the propagated errors of the relevant effective temperatures and steady state densities.
@@
Next, we conduct a series of simulations to investigate the spectrum of driving frequencies wherein equilibrium conditions for reaching a steady state apply. Given that the GFDR holds for $\tau_{s}<\tau_{r}$, we examine our findings through a plot depicting the lower driving frequency, $f_{\text{min}}$, of the two interacting systems. Specifically, we analyze both systems’ effective chemical potential, temperature, and ratio (Fig. 5). We find that as $f_{\text{min}}$ increases, the values of $\mu_{\text{eff}}$ and $T_{\text{eff}}$ change less with the driving frequency. Moreover, the values of $\mu_{\text{eff}}/T_{\text{eff}}$ on both sides agree within measurement error at steady-state only when both driving frequencies are sufficiently high. At these frequencies the driven colloidal suspensions behave as equilibrium colloidal suspensions at elevated temperatures.